OpenStudy (anonymous):

Write the function in vertex form and identify its vertex. h(x) = 3x2 - 24x + 53

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

-(-24x/3x2)+53=0

OpenStudy (anonymous):

in your case you get 4 and \(h(4)=5\) so it is \(h(x)=3(x-4)^2+5\)

OpenStudy (anonymous):

a. h(x) = 3(x - 4)2 + 5; (5, 4) b. h(x) = 3(x - 4)2 + 5; (4, 5)

OpenStudy (anonymous):

wud it be b.???